8. value 10.00 points Stone Sour Co. has an ROA of 11 percent and a payout ratio of 23 percent. What is its internal growth rate? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Internal growth rate Hints References eBook&Resources Hint #1 Check my work
Answer #1

Solution-

Return on assets (ROA) = 11%

Dividend Payout ratio = 23%

Retention ratio = 1- Dividend payout ratio

Retention ratio = 1-0.23

Retention ratio= 0.77

Internal growth rate = (Return on assets * Retention ratio)/ [ 1- ( Return on assets* Retention ratio)]

Internal growth rate = (0.11*0.77)/[1-(0.11*0.77)]

Internal growth rate=0.0847/(1-0.0847)

Internal growth rate = 0.0847/0.9153

Internal growth rate = 0.09253

Internal growth rate = 9.25%